# Characterisation of the big Ramsey degrees of the generic partial order

### v4

@@ Abstract.
As a result of 33 intercontinental Zoom calls, we characterise big Ramsey degrees of the generic partial order. This is an infinitary extension of the well known fact that finite partial orders endowed with linear extensions form a Ramsey class (this result was announced by Nešetřil and Rödl in 1984 with first published proof by Paoli, Trotter and Walker in 1985). Towards this, we refine earlier upper bounds obtained by Hubička based on a new connection of big Ramsey degrees to the Carlson–Simpson theorem and we also introduce a new technique of giving lower bounds using an iterated application of the upper-bound theorem.
@@ 1. Introduction
We characterise the big Ramsey degrees of (finite substructures of) $\mathbf{P}$ in a similar fashion as the authors did in [3] for binary finitely constrained free amalgamation classes. Similarly to that case, our characterisation is in terms of a tree-like object we call a “poset diary” where each level has exactly one critical event. It follows that any poset diary which codes the generic poset is a big Ramsey structure for $\mathbf{P}$ (Definition 1.3 of [43]).
Relative to the small Ramsey degree case (see e.g. [36, 9, 21]), there are fewer classes of structures for which big Ramsey degrees are fully understood. The following is the current state of the art:
@@
    In 1979, Devlin refined upper bounds by Laver and characterised big Ramsey degrees of the rational linear order [16, 41].
@@
    In 2006 Laflamme, Sauer and Vuksanović characterised big Ramsey degrees of the Rado (or random) graph and related random structures in binary languages [27]. Actual numbers were counted by Larson [28].
@@
    In 2008 Nguyen Van Thé characterised big Ramsey degrees of the ultrametric Urysohn spaces [35].
@@
    In 2010 Laflamme, Nguyen Van Thé and Sauer [26] characterised the big Ramsey degrees of the dense local order, $\mathbf{S}(2)$ and the $\mathbb{Q}_{n}$ structures.
@@
    In 2020 Coulson, Dobrinen, and Patel in  [13] and  [14] characterised the big Ramsey degrees of homogeneous binary relational structures which satisfy SDAP+, recovering work in [27] and [26] and characterising the big Ramsey degrees of the ordered versions of structures in [27], the generic $n$-partite and generic ordered $n$-partite graphs, and the $(\mathbb{Q}_{\mathbb{Q}})_{n}$ hierarchy.
@@
    In 2020 Barbosa characterised the big Ramsey degrees of the directed circular orders $\mathbf{S}(k)$, $k\geq 2$,  [8].
@@
    In 2020 a characterisation of the big Ramsey degrees of the triangle-free Henson graph was obtained by Dobrinen [18] and independently by the remaining authors of this article. Joining efforts, the authors were able to fully characterise big Ramsey degrees of free amalgamation classes in finite binary languages described by finitely many forbidden cliques [3].
If one draws an analogy to the small Ramsey results, many of the aforementioned characterisations can be seen as infinitary generalisations of special cases of the Nešetřil–Rödl theorem [32]. Small Ramsey degrees (or the Ramsey expansions satisfying the expansion property, see [36]) are known for many other classes including the class of partial orders with linear extensions [33, 37, 39, 30, 40, 34], metric spaces [31, 15, 30] and other examples satisfying rather general structural conditions [22] (see also recent surveys [21, 25]).
In this paper, we characterize the big Ramsey degrees of the generic partial order. This class represents an important new example since its finitary counterpart is not a consequence of the Nešetřil–Rödl theorem [32], nor is it of the form covered by [14]. To complete our characterisation, we further refine the proof techniques introduced in [20] used to prove finite upper bounds, which were based on the Carlson–Simpson theorem, and we also give a fundamentally different approach than in [3] to prove the lower bounds. The techniques introduced in this paper generalize to other classes as we briefly outline in the final section. However, to keep the paper simple, we did not attempt to state the results in the greatest possible generality.
@@
[…]
Theorem 1.1 (Hubička [20]).
@@
Our construction makes use of the following partial order introduced in [20]; this is closely related to the tree of 1-types within a fixed enumeration of $\mathbf{P}$ discussed in Section 3. Let
@@
    $\begin{split}\qquad\overline{S}_{\ell+1}&=\{z\in\overline{S}_{\ell}:z\mathbin{<_{\mathrm{lex}}}v\hbox{ and }z\perp v\}^{\frown}\mathrm{X}\\ &\qquad\cup\{z\in\overline{S}_{\ell}:z\mathbin{<_{\mathrm{lex}}}v\hbox{ and }z\mathbin{\not\perp}v\}^{\frown}\mathrm{L}\\ &\qquad\cup\{v^{\frown}\mathrm{L}\}\\ &\qquad\cup\{z\in\overline{S}_{\ell}:v\mathbin{<_{\mathrm{lex}}}z\mathbin{<_{\mathrm{lex}}}w\}^{\frown}\mathrm{X}\\ &\qquad\cup\{{w}^{\frown}\mathrm{R}\}\\ &\qquad\cup\{z\in\overline{S}_{\ell}:w\mathbin{<_{\mathrm{lex}}}z\hbox{ and }w\perp z\}^{\frown}\mathrm{X}\\ &\qquad\cup\{z\in\overline{S}_{\ell}:w\mathbin{<_{\mathrm{lex}}}z\hbox{ and }w\not\perp z\}^{\frown}\mathrm{R}.\end{split}$
@@
[…]
$T(\mathbf{A}_{2})$ =$=\{\{\mathrm{X}\mathrm{R},\mathrm{R}\mathrm{X}\mathrm{X}\},\{\mathrm{X}\mathrm{R}\mathrm{X},\mathrm{R}\mathrm{X}\}\},$
$T(\mathbf{C}_{2})$ =$=\{\{\mathrm{X}\mathrm{L},\mathrm{R}\mathrm{R}\mathrm{X}\},\{\mathrm{X}\mathrm{L}\mathrm{X},\mathrm{R}\mathrm{R}\}\}.$
@@
As our main result, we determine the big Ramsey degrees of $\mathbf{P}$ and show that $\mathbf{P}$ admits a big Ramsey structure; while we refer to [43] for the precise definition (see also [23]), a big Ramsey structure for $\mathbf{P}$ is an expansion $\mathbf{P}^{*}$ of $\mathbf{P}$ which encodes the exact big Ramsey degrees for all finite substructures simultaneously in a coherent fashion.
@@
Note that the number of poset-diaries is multiplied by the size of the automorphism group since we define big Ramsey degrees with respect to embeddings (as done, for example, in [43]). Big Ramsey degrees are often defined with respect to substructures (see, for example, [16, 27, 28, 35, 26]) and in that case the degree would be $|T(\mathbf{Q})|$.
@@ 2.2. Trees
For us, a tree is a (possibly empty) partially ordered set $(T,<_{T})$ such that, for every $t\in T$, the set $\{s\in T:s<_{T}t\}$ is finite and linearly ordered by $<_{T}$. All nonempty trees we consider are rooted, that is, they have a unique minimal element called the root of the tree. An element $t\in T$ of a tree $T$ is called a node of $T$ and its level, denoted by $\left|t\right|_{T}$, is the size of the set $\{s\in T:s<_{T}t\}$. Note that the root has level 0. Given a tree $T$ and nodes $s,t\in T$, we say that $s$ is a successor of $t$ in $T$ if $t\mathbin{\leq_{T}}s$. A subtree of a tree $T$ is a subset $T^{\prime}$ of $T$ viewed as a tree equipped with the induced partial ordering.
@@ 2.3. Parameter words
We apply the following Ramsey theorem for parameter words, which is an easy consequence of the Carlson–Simpson theorem[11], see also [41, 24]:
@@ 3. Tree of 1-types
Poset-diaries, which can be compared to Devlin embedding types (see Chapter 6.3 of [41]) or diaries introduced in [3], have an intuitive meaning when understood in the context of the tree of 1-types of $\mathbf{P}$. We now introduce this tree and its enrichment to an aged coding tree and discuss how poset-diaries can be obtained as a suitable abstraction of the aged coding tree.
An enumerated structure is simply a structure $\mathbf{A}$ with underlying set $A=|A|$. Fix a countably infinite enumerated structure $\mathbf{A}$. Given vertices $u,v$ and an integer $n$ satisfying $\min(u,v)\geq n\geq 0$, we write $u\sim^{\mathbf{A}}_{n}v$ and say that $u$ and $v$ have the same (quantifier-free) type over $\{0,1,\ldots,n-1\}$, if the structures induced by $\mathbf{A}$ on $\{0,1,\ldots,n-1,u\}$ and $\{0,1,\ldots,n-1,v\}$ are isomorphic via the map which is the identity on $\{0,...,n-1\}$ and sends $u$ to $v$. We write $[u]^{\mathbf{A}}_{n}$ for the $\sim^{\mathbf{A}}_{n}$-equivalence class of the vertex $u$.
@@
Let $\mathbf{A}$ be a countably infinite (relational) enumerated structure. Given $n<\omega$, write $\mathbb{T}_{\mathbf{A}}(n)=\omega/\!\sim^{\mathbf{A}}_{n}$. A (quantifier-free) 1-type is any member of the disjoint union $\mathbb{T}_{\mathbf{A}}:=\bigsqcup_{n<\omega}\mathbb{T}_{\mathbf{A}}(n)$. We turn $\mathbb{T}_{\mathbf{A}}$ into a tree as follows. Given $x\in\mathbb{T}_{\mathbf{A}}(m)$ and $y\in\mathbb{T}_{\mathbf{A}}(n)$, we declare that $x\leq^{\mathbb{T}}_{\mathbf{A}}y$ if and only if $m\leq n$ and $x\supseteq y$. We will denote by $[n]^{\mathbf{A}}_{n}$ the equivalence class of $n$ under $\sim^{\mathbf{A}}_{n}$.
In the case that we have a Fraïssé class $\cal{K}$ in mind (which for us will always be the class of finite partial orders), we can extend the definition to a finite enumerated $\mathbf{A}\in\cal{K}$ as follows. Fix an enumerated Fraïssé limit $\mathbf{K}$ of $\cal{K}$ which has $\mathbf{A}$ as an initial segment. We then set $\mathbb{T}_{\mathbf{A}}=\mathbb{T}_{\mathbf{K}}({<}|\mathbf{A}|)$. This does not depend on the choice of $\mathbf{K}$.
@@
The tree of 1-types of a given structure is useful for constructing unavoidable colourings as well as for proving upper bounds on big Ramsey degrees; see for instance [27, 17, 19, 13, 14, 44, 20, 6, 7, 3]. We therefore fix an enumerated generic partial order $\mathbf{P}$, and we put $(\mathbb{T},\leq_{\mathbb{T}}\nobreak)=(\mathbb{T}_{\mathbf{P}},\leq^{\mathbb{T}}_{\mathbf{P}})$ and $\mathrm{CT}=\mathrm{CT}^{\mathbf{P}}$. Noting that $\mathbf{P}$ is a binary relational structure and that there are exactly $3$ enumerated posets of size $2$, we have $k=3$; by identifying the symbols $\{\mathrm{L},\mathrm{X},\mathrm{R}\}$ with $\{0,1,2\}$, we can identify $\mathrm{CT}$ as a subtree of $(\Sigma^{*},\sqsubseteq)$. More concretely, given $x\in\mathbb{T}(m)$ and $j<m$, we have:
$\sigma(x)_{j}=\begin{cases}\mathrm{L}&\hbox{if $a\mathbin{<_{\mathbf{P}}}j$ for every (some) $a\in x$,}\\ \mathrm{R}&\hbox{if $j\mathbin{<_{\mathbf{P}}}a$ for every (some) $a\in x$,}\\ \mathrm{X}&\hbox{otherwise.}\end{cases}$
@@
    if there exists $\ell^{\prime}<\min(|u|,|v|)$ such that $(u_{\ell^{\prime}},v_{\ell^{\prime}})=(\mathrm{L},\mathrm{R})$, then for every ′′$\ell^{\prime\prime}<\min(|u|,|v|)$ it holds that $u_{\ell^{\prime\prime}}\mathbin{\leq_{\mathrm{\Sigma}}}v_{\ell^{\prime\prime}}$.
@@
In [44], using ideas implicit in the parallel $1$’s of [17, 38] and pre-$a$-cliques of [19], levels of coding trees are endowed with the structure of aged sets. This means that for every $m<\omega$, every set $S\subseteq\mathrm{CT}(m)$ is equipped with a class of finite $|S|$-labeled structures describing exactly which finite structures can be coded by coding nodes above the members of $S$. For the generic partial order, it will be useful to encode this information slightly differently than in [44, 3], in particular since we want to do this on all of $\Sigma^{*}$, not just on $\mathrm{CT}$.
@@
Given $\ell\geq 0$ and $S\subseteq\Sigma^{*}_{\ell}$, the level structure of $S$ is the structure $\mathbf{S}=(S,\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)$.
@@
For every $\ell>0$ and $S\subseteq\Sigma^{*}_{\ell}$, the level structure $\mathbf{S}=(S,\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)$ satisfies the following properties:
@@
One can, perhaps surprisingly, prove that the class $\mathcal{K}$ of all finite structures satisfying properties (P1), (P2), …, (P7) is an amalgamation class. As a consequence of the construction from Section 5 one gets that for each structure $\mathbf{A}\in\cal{K}$ there exists $\ell>0$ and $S\subseteq\Sigma^{*}_{\ell}$ such that $\mathbf{S}=(S,\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)$ is isomorphic to $\mathbf{A}$.
@@
This interesting phenomenon of constructing a class of approximations (or, using the terminology of [44, 3], the class of all possible aged sets that can appear on some level of the coding tree) of a given amalgamation class exists in other cases. For binary free amalgamation classes, this approximation class corresponds to the union $\bigcup_{\rho}P(\rho)$, where the union is taken over all possible sorts $\rho$ (see [3] for the definitions). However, the theory of aged coding trees and the sets $P(\rho)$ can be defined for any strong amalgamation class in a finite binary language. Note that while for free amalgamation classes the set $P(\rho)$ is always closed under intersections, this need not be the case in general (indeed, it fails for posets).
Another key difference between the free amalgamation case and partial orders is that for free amalgamation classes, we can enumerate the Fraïssé limit in such a way so that going up and left (that is, by a non-relation) in the coding tree is a safe move, i.e., is an embedding of the level structure from one level to another. Indeed, if this were true for the generic partial order and the coding tree $\mathrm{CT}$ we fixed earlier, one could prove upper bounds for the big Ramsey degrees using forcing arguments much as is done for the free amalgamation case in [44]. However, while a weakening of the idea of a “safe direction” does hold for partial orders (see Proposition 6.7), the proof of Lemma 3.4 from [44] breaks in the setting of the generic partial order. However, it is possible that the coding tree Milliken theorem still holds. Below, $\mathrm{AEmb}(\mathrm{CT}^{\mathbf{A}},\mathrm{CT})$ refers to the set of aged embeddings of the coding tree $\mathrm{CT}^{\mathbf{A}}$ into $\mathrm{CT}$, the strong similarity maps that respect coding nodes and level structures (see Definition 2.3 of [44]).

### v5

@@ Abstract.
As a result of 33 intercontinental Zoom calls, we characterise big Ramsey degrees of the generic partial order. This is an infinitary extension of the well known fact that finite partial orders endowed with linear extensions form a Ramsey class (this result was announced by Nešetřil and Rödl in 1984). Towards this, we refine earlier upper bounds obtained by Hubička based on a new connection of big Ramsey degrees to the Carlson–Simpson theorem and we also introduce a new technique of giving lower bounds using an iterated application of the upper-bound theorem.
@@ 1. Introduction
We characterise the big Ramsey degrees of (finite substructures of) $\mathbf{P}$ in a similar fashion as the authors did in [BCD+24] for binary finitely constrained free amalgamation classes. Similarly to that case, our characterisation is in terms of a tree-like object we call a “poset diary” where each level has exactly one critical event. It follows that any poset diary which codes the generic poset is a big Ramsey structure for $\mathbf{P}$ (Definition 1.3 of [Zuc19]).
Relative to the small Ramsey degree case (see e.g. [HK26, NVT15, Bod15, Hub20]), there are fewer classes of structures for which big Ramsey degrees are fully understood. The following is the current state of the art (see [HZ25, HK26, Dob23c]):
@@
    In 1979, Devlin refined upper bounds by Laver and characterised big Ramsey degrees of the rational linear order [Dev79, Tod10], see also Sequence A000182 of OEIS [OEI25].
@@
    In 2006 Laflamme, Sauer and Vuksanović characterised big Ramsey degrees of the Rado (or random) graph and related random structures in binary languages [LSV06]. Actual numbers were counted by Larson [Lar08], see also Sequence A293158 of OEIS [OEI25].
@@
    In 2008 Nguyen Van Thé characterised big Ramsey degrees of the ultrametric Urysohn spaces [NVT08].
@@
    In 2010 Laflamme, Nguyen Van Thé and Sauer [LNVTS10] characterised the big Ramsey degrees of the dense local order, $\mathbf{S}(2)$ and the $\mathbb{Q}_{n}$ structures.
@@
    In 2020 Coulson, Dobrinen, and Patel in  [CDP22a] and  [CDP22b] characterised the big Ramsey degrees of homogeneous binary relational structures which satisfy SDAP+, recovering work in [LSV06] and [LNVTS10] and characterising the big Ramsey degrees of the ordered versions of structures in [LSV06], the generic $n$-partite and generic ordered $n$-partite graphs, and the $(\mathbb{Q}_{\mathbb{Q}})_{n}$ hierarchy.
@@
    In 2020 Barbosa characterised the big Ramsey degrees of the directed circular orders $\mathbf{S}(k)$, $k\geq 2$, [Bar20].
@@
    In 2020 a characterisation of the big Ramsey degrees of the triangle-free Henson graph was obtained by Dobrinen [Dob20b] and independently by the remaining authors of this article. Joining efforts, the authors were able to fully characterise big Ramsey degrees of free amalgamation classes in finite binary languages described by finitely many forbidden cliques [BCD+24]. Exact counts for the case of $K_{3}$ and $K_{4}$-free graphs were recently determined by Hubička, Konečný, Vodseďálek, and Zucker [HKVZ25, Vod25], see also Sequence A387346 of OEIS [OEI25].
If one draws an analogy to the small Ramsey results, many of the aforementioned characterisations can be seen as infinitary generalisations of special cases of the Nešetřil–Rödl theorem [NR77]. Small Ramsey degrees (or the Ramsey expansions satisfying the expansion property, see [NVT15]) are known for many other classes including the class of partial orders with linear extensions [NR84, PTW85, Sok12, Maš18, SZ17, NR18], metric spaces [Neš07, DR12, Maš18] and other examples satisfying rather general structural conditions [HN19] (see also recent surveys [HK26, Hub20, Kon23]).
In this paper, we characterize the big Ramsey degrees of the generic partial order. This class represents an important new example since its finitary counterpart is not a consequence of the Nešetřil–Rödl theorem [NR77], nor is it of the form covered by [CDP22b]. To complete our characterisation, we further refine the proof techniques introduced in [Hub25] used to prove finite upper bounds, which were based on the Carlson–Simpson theorem, and we also give a fundamentally different approach than in [BCD+24] to prove the lower bounds. The techniques introduced in this paper generalize to other classes as we briefly outline in the final section. However, to keep the paper simple, we did not attempt to state the results in the greatest possible generality.
@@
For a countably infinite structure $\mathbf{B}$ and a finite substructure $\mathbf{A}$, the big Ramsey degree of $\mathbf{A}$ in $\mathbf{B}$ is the least number $\ell\in\mathbb{N}\cup\{\infty\}$ such that $\mathbf{B}\longrightarrow(\mathbf{B})^{\mathbf{A}}_{k,\ell}$ for every $k\in\mathbb{N}$. Similarly, if $\mathcal{K}$ is a class of finite structures and $\mathbf{A}\in\mathcal{K}$, the small Ramsey degree of $\mathbf{A}$ in $\mathcal{K}$ is the least number $\ell\in\mathbb{N}\cup\{\infty\}$ such that for every $\mathbf{B}\in\mathcal{K}$ and $k\in\mathbb{N}$ there exists $\mathbf{C}\in\mathcal{K}$ such that $\mathbf{C}\longrightarrow(\mathbf{B})^{\mathbf{A}}_{k,l}$.
A structure is homogeneous if every isomorphism between two of its finite substructures extends to an automorphism. It is well known that up to isomorphism there is a unique homogeneous partial order $\mathbf{P}=(P,\leq_{\mathbf{P}})$ such that every countable partial order has an embedding into $\mathbf{P}$. The order $\mathbf{P}$ is called the generic partial order (see e.g. [Mac11]). We refine the following recent result.
Theorem 1.1 (Hubička [Hub25]).
@@
Our construction makes use of the following partial order introduced in [Hub25]; this is closely related to the tree of 1-types within a fixed enumeration of $\mathbf{P}$ discussed in Section 3. Let
@@
[…]
@@
$T(\mathbf{A}_{1})=T(\mathbf{C}_{1})$ $=\{\emptyset\},$
$T(\mathbf{A}_{2})$ $=\{\{\mathrm{X}\mathrm{R},\mathrm{R}\mathrm{X}\mathrm{X}\},\{\mathrm{X}\mathrm{R}\mathrm{X},\mathrm{R}\mathrm{X}\}\},$
$T(\mathbf{C}_{2})$ $=\{\{\mathrm{X}\mathrm{L},\mathrm{R}\mathrm{R}\mathrm{X}\},\{\mathrm{X}\mathrm{L}\mathrm{X},\mathrm{R}\mathrm{R}\}\}.$
@@
As our main result, we determine the big Ramsey degrees of $\mathbf{P}$ and show that $\mathbf{P}$ admits a big Ramsey structure; while we refer to [Zuc19] for the precise definition (see also [HZ25]), a big Ramsey structure for $\mathbf{P}$ is an expansion $\mathbf{P}^{*}$ of $\mathbf{P}$ which encodes the exact big Ramsey degrees for all finite substructures simultaneously in a coherent fashion.
@@
Note that the number of poset-diaries is multiplied by the size of the automorphism group since we define big Ramsey degrees with respect to embeddings (as done, for example, in [Zuc19]). Big Ramsey degrees are often defined with respect to substructures (see, for example, [Dev79, LSV06, Lar08, NVT08, LNVTS10]) and in that case the degree would be $|T(\mathbf{Q})|$.
@@ 2.2. Trees
For us, a tree is a (possibly empty) partially ordered set $(T,<_{T})$ such that, for every $t\in T$, the set $\{s\in T:s<_{T}t\}$ is finite and linearly ordered by $_{T}$. All nonempty trees we consider are rooted, that is, they have a unique minimal element called the root of the tree. An element $t\in T$ of a tree $T$ is called a node of $T$ and its level, denoted by $\left|t\right|_{T}$, is the size of the set $\{s\in T:s<_{T}t\}$. Note that the root has level 0. Given a tree $T$ and nodes $s,t\in T$, we say that $s$ is a successor of $t$ in $T$ if $t\mathbin{\leq_{T}}s$. A subtree of a tree $T$ is a subset $T^{\prime}$ of $T$ viewed as a tree equipped with the induced partial ordering.
@@ 2.3. Parameter words
We apply the following Ramsey theorem for parameter words, which is an easy consequence of the Carlson–Simpson theorem[CS84], see also [Tod10, Kar13]:
@@ 3. Tree of 1-types
Poset-diaries, which can be compared to Devlin embedding types (see Chapter 6.3 of [Tod10]) or diaries introduced in [BCD+24], have an intuitive meaning when understood in the context of the tree of 1-types of $\mathbf{P}$. We now introduce this tree and its enrichment to an aged coding tree and discuss how poset-diaries can be obtained as a suitable abstraction of the aged coding tree.
An enumerated structure is simply a structure $\mathbf{A}$ with underlying set $A=|A|$. Fix a countably infinite enumerated structure $\mathbf{A}$. Given vertices $u,v$ and an integer $n$ satisfying $\min(u,v)\geq n\geq 0$, we write $u\sim^{\mathbf{A}}_{n}v$ and say that $u$ and $v$ have the same (quantifier-free) type over $\{0,1,\ldots,n-1\}$, if the structures induced by $\mathbf{A}$ on $\{0,1,\ldots,n-1,u\}$ and $\{0,1,\ldots,n-1,v\}$ are isomorphic via the map which is the identity on $\{0,...,n-1\}$ and sends $u$ to $v$. We write $u]^{\mathbf{A}}_{n}$ for the $\sim^{\mathbf{A}}_{n}$-equivalence class of the vertex $u$.
@@
Let $\mathbf{A}$ be a countably infinite (relational) enumerated structure. Given $n<\omega$, write $\mathbb{T}_{\mathbf{A}}(n)=\omega/\!\sim^{\mathbf{A}}_{n}$. A (quantifier-free) 1-type is any member of the disjoint union $\mathbb{T}_{\mathbf{A}}:=\bigsqcup_{n<\omega}\mathbb{T}_{\mathbf{A}}(n)$. We turn $\mathbb{T}_{\mathbf{A}}$ into a tree as follows. Given $x\in\mathbb{T}_{\mathbf{A}}(m)$ and $y\in\mathbb{T}_{\mathbf{A}}(n)$, we declare that $x\leq^{\mathbb{T}}_{\mathbf{A}}y$ if and only if $m\leq n$ and $x\supseteq y$. We will denote by $n]^{\mathbf{A}}_{n}$ the equivalence class of $n$ under $\sim^{\mathbf{A}}_{n}$.
In the case that we have a Fraïssé class $\mathcal{K}$ in mind (which for us will always be the class of finite partial orders), we can extend the definition to a finite enumerated $\mathbf{A}\in\mathcal{K}$ as follows. Fix an enumerated Fraïssé limit $\mathbf{K}$ of $\mathcal{K}$ which has $\mathbf{A}$ as an initial segment. We then set $\mathbb{T}_{\mathbf{A}}=\mathbb{T}_{\mathbf{K}}({<}|\mathbf{A}|)$. This does not depend on the choice of $\mathbf{K}$.
@@
The tree of 1-types of a given structure is useful for constructing unavoidable colourings as well as for proving upper bounds on big Ramsey degrees; see for instance [LSV06, Dob20a, Dob23b, CDP22a, CDP22b, Zuc22, Hub25, BCH+19, BCH+22, BCD+24, Dob20c, Dob23a, HZ25, HK26]. We therefore fix an enumerated generic partial order $\mathbf{P}$, and we put $(\mathbb{T},\leq_{\mathbb{T}}\penalty 10000)=(\mathbb{T}_{\mathbf{P}},\leq^{\mathbb{T}}_{\mathbf{P}})$ and $\mathrm{CT}=\mathrm{CT}^{\mathbf{P}}$. Noting that $\mathbf{P}$ is a binary relational structure and that there are exactly $3$ enumerated posets of size $2$, we have $k=3$; by identifying the symbols $\{\mathrm{L},\mathrm{X},\mathrm{R}\}$ with $\{0,1,2\}$, we can identify $\mathrm{CT}$ as a subtree of $(\Sigma^{*},\sqsubseteq)$. More concretely, given $x\in\mathbb{T}(m)$ and $j<m$, we have:
[…]
@@
    if there exists $\ell^{\prime}<\min(|u|,|v|)$ such that $(u_{\ell^{\prime}},v_{\ell^{\prime}})=(\mathrm{L},\mathrm{R})$, then for every $\ell^{\prime\prime}<\min(|u|,|v|)$ it holds that $u_{\ell^{\prime\prime}}\mathbin{\leq_{\mathrm{\Sigma}}}v_{\ell^{\prime\prime}}$.
@@
In [Zuc22], using ideas implicit in the parallel $1$’s of [Dob20a, Sau98] and pre-$a$-cliques of [Dob23b], levels of coding trees are endowed with the structure of aged sets. This means that for every $m<\omega$, every set $S\subseteq\mathrm{CT}(m)$ is equipped with a class of finite $|S|$-labeled structures describing exactly which finite structures can be coded by coding nodes above the members of $S$. For the generic partial order, it will be useful to encode this information slightly differently than in [Zuc22, BCD+24], in particular since we want to do this on all of $\Sigma^{*}$, not just on $\mathrm{CT}$.
@@
Given $\ell\geq 0$ and $S\subseteq\Sigma^{*}_{\ell}$, the level structure of $S$ is the structure $\mathbf{S}=(S,\allowbreak\mathbin{\leq_{\mathrm{lex}}}\penalty 10000,\allowbreak\preceq\penalty 10000,\allowbreak\trianglelefteq)$.
@@
For every $\ell>0$ and $S\subseteq\Sigma^{*}_{\ell}$, the level structure $\mathbf{S}=(S,\allowbreak\mathbin{\leq_{\mathrm{lex}}}\penalty 10000,\allowbreak\preceq\penalty 10000,\allowbreak\trianglelefteq)$ satisfies the following properties:
@@
One can, perhaps surprisingly, prove that the class $\mathcal{K}$ of all finite structures satisfying properties (P1), (P2), …, (P7) is an amalgamation class. As a consequence of the construction from Section 5 one gets that for each structure $\mathbf{A}\in\mathcal{K}$ there exists $\ell>0$ and $S\subseteq\Sigma^{*}_{\ell}$ such that $\mathbf{S}=(S,\allowbreak\mathbin{\leq_{\mathrm{lex}}}\penalty 10000,\allowbreak\preceq\penalty 10000,\allowbreak\trianglelefteq)$ is isomorphic to $\mathbf{A}$.
@@
This interesting phenomenon of constructing a class of approximations (or, using the terminology of [Zuc22, BCD+24], the class of all possible aged sets that can appear on some level of the coding tree) of a given amalgamation class exists in other cases. For binary free amalgamation classes, this approximation class corresponds to the union $\bigcup_{\rho}P(\rho)$, where the union is taken over all possible sorts $\rho$ (see [BCD+24] for the definitions). However, the theory of aged coding trees and the sets $P(\rho)$ can be defined for any strong amalgamation class in a finite binary language. Note that while for free amalgamation classes the set $P(\rho)$ is always closed under intersections, this need not be the case in general (indeed, it fails for posets).
Another key difference between the free amalgamation case and partial orders is that for free amalgamation classes, we can enumerate the Fraïssé limit in such a way so that going up and left (that is, by a non-relation) in the coding tree is a safe move, i.e., is an embedding of the level structure from one level to another. Indeed, if this were true for the generic partial order and the coding tree $\mathrm{CT}$ we fixed earlier, one could prove upper bounds for the big Ramsey degrees using forcing arguments much as is done for the free amalgamation case in [Zuc22]. However, while a weakening of the idea of a “safe direction” does hold for partial orders (see Proposition 6.7), the proof of Lemma 3.4 from [Zuc22] breaks in the setting of the generic partial order. However, it is possible that the coding tree Milliken theorem still holds. Below, $\mathrm{AEmb}(\mathrm{CT}^{\mathbf{A}},\mathrm{CT})$ refers to the set of aged embeddings of the coding tree $\mathrm{CT}^{\mathbf{A}}$ into $\mathrm{CT}$, the strong similarity maps that respect coding nodes and level structures (see Definition 2.3 of [Zuc22]).